# Expert-elicitation method for non-parametric joint priors using normalizing flows

### v1

@@ 1 Introduction
A whole field of research, commonly referred to as (expert) prior elicitation, has emerged around the question of how to gather expert knowledge and translate it into appropriate prior distributions [69, 50, 23]. Garthwaite et al [28] identified four key stages in a prior elicitation process:
@@
    Setup stage: In this stage, the problem is defined, an expert is selected, and the quantities to be elicited from the expert (referred to as target quantities) are determined;
@@
In this context, elicitation methods aim to provide a systematic and formal procedure for deriving prior distributions based on expert-elicited statistics. Early elicitation methods primarily tackled the problem by seeking analytical solutions, such as conjugate models or problem-specific transformation functions resulting in highly model-specific prior elicitation methods [see 50, for a recent review]. This model dependence is partly due to the use of direct elicitation techniques, which involve asking experts directly about model parameters [69, 23]. This approach is problematic not only due to its inherent model dependence but also because model parameters are often challenging for domain experts to interpret meaningfully. Consequently, the quality of expert input can be questionable. In response to these limitations, substantial efforts have been made to develop methods that enable more flexible and interpretable prior specification for domain experts.
@@
Building in particular on the work of Hartmann et al [38], da Silva et al [67], Manderson and Goudie [47], we previously introduced a simulation-based framework with a highly modular structure that closely aligns with the elicitation process described by Garthwaite et al [28]. The advantage of this modular structure is its flexibility, allowing for easy adaptation to different types of methods. In our previous work [6], we demonstrated how to use this framework to learn parametric prior distributions for the model parameters. In this paper, we show how the same simulation-based framework can be used to learn flexible (i.e., non-parametric) joint priors for the model parameters with only minor adjustments to the workflow. In both approaches, we employ mini-batch stochastic gradient descent as the optimization method. However, our framework supports different optimization methods thanks to its modular design.
@@ 2 Methodology
In this paper, we propose an elicitation method that builds on and extends our recently introduced simulation-based framework [6]. The overall idea of this framework closely resembles the approach of prior predictive checks [32]: We simulate from the joint model $p(\theta,y)$ and assess how well the resulting prior predictions align with the expert’s expectations. If there is a discrepancy between the expert’s expectations and the simulations, the prior specification needs to be adjusted accordingly. Figure 1 provides a graphical representation of the framework and Table 1 offers a summary of the symbols and notation used in the following sections. The general workflow of our framework can be summarized as follows:
@@ 2.1 Prior Specification & Learning
In our previous work [6], we introduced the simulation-based framework along with a method for learning a set of independent parametric prior distributions, $p(\theta_{k}\mid\lambda_{k})$, for the model parameters $\theta_{k}$ with $k=1,\ldots,K$. In this paper, we extend our framework by introducing a method to learn a joint non-parametric prior distribution $p_{\lambda}(\theta)=p_{\lambda}(\theta_{1},\ldots,\theta_{K})$ over all model parameters $\theta$. In both the parametric and non-parametric approaches, the objective of the optimization process is to learn the hyperparameters $\lambda$ that minimizes the discrepancy between the model simulations and the expert expectations. However, the interpretation of these hyperparameters differs between approaches, which we emphasize by using a different notation for the prior distributions. In the non-parameteric approach, we employ normalizing flows (NFs) to induce a flexible family of prior distributions which entail specialized deep neural networks with trainable parameters $\lambda$ [44]. Thus, in the non-parametric approach, $\lambda$ represents the weights of the deep neural networks within the NFs, whereas in the parametric approach, $\lambda$ denotes the parameters of the associated prior distribution families (i.e., the prior hyperparameters).
@@
To learn a non-parametric joint prior for the model parameters, we use normalizing flows (NFs), a family of deep generative models that offers tractable density evaluation [44]. NFs transform a simple probability distribution, the base distribution $p(u)$, into a complex target distribution — in our case, the joint prior $p_{\lambda}(\theta)$ — through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$ [44]. The composition of invertible functions $g_{\lambda}$ yields an explicit form of the density $p_{\lambda}(\theta)$ via the change of variables formula
@@
Typically, $p(u)$ is chosen to be a simple density, such as a standard Gaussian or a uniform distribution [44]. Figure 2 depicts the use of NFs for learning a non-parameteric prior distribution within our simulation-based framework.
Figure 2: Learning non-parametric prior distributions using normalizing flows within our simulation-based framework. Normalizing flows transform a simple probability distribution, the base distribution $p(u)$, into a complex target distribution — in our case, the joint prior $p_{\lambda}(\theta)$ — through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$.
@@ 2.2 Target Quantities & Elicitation Techniques
In our elicitation method, we represent an elicitation technique as a function $f_{j}$ of a target quantity $z_{i}$. The resulting set of elicited target quantities is referred to as elicited statistics,  $\{t_{m}(\lambda)\}$, where $t_{m}(\lambda)=f_{j}(z_{i})$ and $m$ refers to the corresponding $i\times j$ combination. In the following, we sometimes keep the dependency on $\lambda$ implicit and simply write $t_{m}$ instead of $t_{m}(\lambda)$.
@@ 2.3 Model-implied & Expert-elicited Statistics
After determining the set of target quantities and corresponding elicitation techniques, we can proceed to gather information from the expert through specific prior elicitation protocols [35, 12, 2]. This process may be preceded by a training phase to familiarize the expert with the elicitation tasks [69]. In prior elicitation research, an entire subfield has focused on determining how to conduct optimal expert interviews, taking into account factors such as cognitive biases and heuristics. While it is beyond the scope of this paper to review this literature, several comprehensive reviews are available [e.g., 50, 69, 23]. At this point, we will skip the actual interrogation process and assume that the expert information has been successfully gathered, providing us with a set of expert-elicited statistics, denoted by $\hat{t}_{m}$.
@@
As an example, consider a Binomial regression model with a dummy-coded categorical predictor $x$ with two levels and model parameters $\theta=(\theta_{1},\theta_{2})$ representing the intercept and slope. We decide to query the expert about the quartiles of the expected outcome variable $y$ conditional on each group, $\text{gr}_{1}$ and $\text{gr}_{2}$. That is, we obtain six expert-elicited statistics $Q_{p}(y_{\text{gr}_{i}})$ with $p=0.25,0.50,0.75$ and $i=1,2$. To obtain the corresponding model-implied statistics, we would first sample the model parameters from the initial prior distribution, $\theta^{(s)}\sim p_{\lambda_{0}}(\theta)$ and compute the linear predictor $\eta^{(s)}_{\text{gr}_{i}}=\theta_{1}^{(s)}+\theta_{2}^{(s)}x_{i}$ transformed by an inverse-link function $\sigma$ to obtain the implied event probabilities $p^{(s)}_{\text{gr}_{i}}=\sigma\left(\eta^{(s)}_{\text{gr}_{i}}\right)$. Then we would sample from the likelihood providing us with samples from the implied prior predictive distribution, $y^{(s)}_{\text{gr}_{i}}\sim\text{Binomial}(p^{(s)}_{\text{gr}_{i}},N)$. We repeat this sampling process $S$ times resulting in sets $\{y^{(s)}_{\text{gr}_{i}}\}$ of $S$ samples. Finally, we compute the quartiles from these samples, resulting in the six model-implied statistics $Q_{p}(\{y^{(s)}_{\text{gr}_{i}}\})$ with $p=0.25,0.50,0.75$ and $i=1,2$.
@@ 2.4 Discrepancy Evaluation & Training
where $\alpha_{m}$ denote the weights. The choice of weights $\alpha_{m}$ can be guided by different considerations: One approach is to manually adjust the weights based on the importance of the individual loss components from the expert’s perspective [72]. Alternatively, automated weighting methods, known as loss-balancing methods, can be used to optimally balance the contribution of each individual term to the total gradient [46, 13, 4].
@@ 2.5 Evaluation of Training Results
Once the elicited statistics have been learned accurately, we can examine the learned prior distribution that corresponds to this set of expert-elicited statistics. However, there are often multiple optimal values $\lambda^{*}$ that align closely with the elicited statistics yet result in different prior distributions [67, 47, 69]. This lack of uniqueness is anticipated, given the limited information provided by the set of elicited statistics compared to the complexity of the generative model [47]. Furthermore, when target quantities refer solely to the observable space, we only gain access to the model’s overall uncertainty with limited ability to differentiate between aleatoric and epistemic uncertainty [53, 60]. Aleatoric uncertainty, by definition, is irreducible; however, several methods exist to reduce epistemic uncertainty. These include (i) selecting more informative target quantities, (ii) eliciting more detailed expert information (e.g., increasing the number of quantiles in quantile-based elicitation), (iii) adding theory-informed regularization terms to constrain the search space [as illustrated in 47], (iv) improving the initialization method to achieve better starting points, and (v) employing more advanced optimization algorithms when dealing with a large number of model parameters [53].
@@
Using the above approach allows us to assess the self-consistency of the algorithm, the degree of non-identifiability (i.e., how informative are the elicited statistics), and the suitability of the chosen algorithm parameters used for training NFs with mini-batch SGD. Additionally, we can use this ground truth as a benchmark to assess whether the learned prior distributions approximate it reasonably well, even though non-identifiability may prevent us from exactly recovering the underlying “true” prior distribution.
@@ 3 Implementation
In all simulation studies, we use a batch size of $128$. To simulate from an oracle (discussed in Section 4.1), we draw once $S=10,000$ samples from the true joint prior and then compute the corresponding set of “expert”-elicited statistics. During training with mini-batch SGD we draw $S=200$ samples from the joint prior. To evaluate the discrepancy between the model-implied and expert-elicited statistics, we use a biased Maximum Mean Discrepancy (MMD) loss with an energy kernel [36, 25], following our previous success with the metric [6]. For the correlation between parameters, however, we apply a squared-error loss (for details see Section 4.1). The total loss is computed as a weighted sum of the individual loss components, with unit weights assigned to the MMD loss components and a weight of $0.1$ applied to the squared-error loss components.
The architecture of the NFs is rather simple compared to typical applications in deep learning [e.g., 42] and consists of a standard multivariate Gaussian as a base distribution and three affine coupling blocks. Each coupling block consists of two dense layers with $128$ units and ReLU activation functions. In Simulation Study 1, the number of learnable hyperparameters is $\mid\lambda\mid=202,776$, while in the case studies corresponding to Simulation Study 2, it is $\mid\lambda\mid=205,872$. In Simulation Study 1, we assume a discrete likelihood and use the softmax-gumble trick to compute the gradients of a discrete random variable [for more information, see 6]. The gradients of the total loss with respect to the hyperparameters are used to update the hyperparameters $\lambda$. Optimization was performed with the Adam optimizer for $600$ epochs in Simulation Study 1 and $1500$ epochs in Simulation Study 2. We used a fixed learning rate of $0.00025$ for Simulation Study 2, Scenario 1 & 2, and $0.0001$ for Simulation Study 1 and Stimulation Study 2, Scenario 3. These algorithm (hyper-)parameter values were obtained by manual tuning until good performance across all of our case studies was achieved.
All simulations were implemented in Python 3.11.10 using the following libraries: TensorFlow 2.14 [1], TensorFlow Probability 0.22.1 [18], NumPy 1.24.3 [37], pandas 2.2.3 [49], and BayesFlow 1.1.6 [63]. Simulations were run on the Linux HPC cluster at Technical University Dortmund (LiDO3) on high-end CPUs.
@@ 4.1 General Setup
[…]
@@
Table 2: Overview of generative models and ground truth in the simulation studies. Symbols: $\mathbf{R}$ denotes the correlation matrix and $s$ the vector of standard deviations.
@@
For the simulation studies in this paper, we simulate expert input using synthetic data generated from a predefined joint prior representing the ground truth (i.e., oracle, see Section 2.5 for details). For simplicity, we use a parametric prior with fixed hyperparameter values as ground truth for each case study. Further details on the exact specification of the “true” joint prior is provided in the respective case-study sections and in Table 2.
@@
Following recommendations in the field, we focused on observable quantities when selecting target quantities for our simulation studies (see Section 2.2 for details). For the Binomial regression model, we used the expected observations of the outcome variable conditional on two observations of the continuous predictor, $y\mid x_{0}$ and $y\mid x_{1}$. For the normal regression model, we considered the prior predictive observations conditional on the three groups, $y\mid\text{gr}_{i}$ for $i=1,2,3$.
@@
Regarding the selection of elicitation techniques, we follow the recommendation to query quantiles from a domain expert [40]. For each target quantity, except for the correlation information, we elicit five quantiles: $5\%,25\%,50\%,75\%,95\%$. For the correlation values, we assume that the expert provides a single point estimate corresponding to a moment-based elicitation approach. To assess whether the selected set of elicited statistics is sufficiently informative to determine a prior distribution for the model parameters, we conduct a sensitivity analysis for each case study.
@@
We run each simulation study 30 times with different random seeds. For each replication, we calculate the slope of the loss trajectory over the last 100 epochs and perform a visual convergence check for the five replications with the highest absolute slopes (i.e., worst cases). In this visual check, we examine the trajectories of the total loss, the loss components related to the elicited statistics, and the mean and standard deviation for each marginal prior across epochs. If this visual inspection indicates satisfactory convergence, we consider the convergence of all other replications to be successful as well. Additionally, we check the match between the final model-implied and expert-elicited statistics. Finally, we examine the learned prior distributions for each replication and compute the model average across the 30 learned prior distributions.
Summary of Analysis Workflow
Table 3 provides a summary of the analysis workflow followed in each of the subsequent simulation studies.

### v2

@@ 1 Introduction
A distinct field of research, commonly known as (expert) prior elicitation, has developed around the question of how to gather expert knowledge and translate it into appropriate prior distributions. This area of study has a long history, dating back to the 1960s [75, 43, 42], and continues to be an active area of research today [71, 52, 24]. Garthwaite et al [30] identified four key stages in a prior elicitation process:
@@
    Setup stage: In this stage, the problem is defined, an expert is selected, and the quantities to be elicited from the expert (in this paper referred to as target quantities) are determined;
@@
In this context, elicitation methods aim to provide a systematic and formal procedure for deriving prior distributions based on expert-elicited statistics. Early elicitation methods primarily tackled the problem by seeking analytical solutions, such as conjugate models or problem-specific transformation functions resulting in highly model-specific prior elicitation methods [see 52, for a recent review]. This model dependence is partly due to the use of direct elicitation techniques, which involve asking experts directly about model parameters [71, 24]. This approach is problematic not only due to its inherent model dependence but also because model parameters are often challenging for domain experts to interpret meaningfully. Consequently, the quality of expert input can be questionable. In response to these limitations, substantial efforts have been made to develop methods that enable more flexible and interpretable prior specification for domain experts.
@@
Building in particular on the work of Hartmann et al [40], da Silva et al [69], Manderson and Goudie [49], we previously introduced a simulation-based framework with a highly modular structure that closely aligns with the elicitation process described by Garthwaite et al [30]. The advantage of this modular structure is its flexibility, allowing for easy adaptation to different types of methods. In our previous work [6], we demonstrated how to use this framework to learn parametric prior distributions for the model parameters. In this paper, we show how the same simulation-based framework can be used to learn flexible (i.e., non-parametric) joint priors for the model parameters with only minor adjustments to the workflow. In both approaches, we employ mini-batch stochastic gradient descent as the optimization method.
@@ 2 Methodology
In this paper, we propose an elicitation method that extends our recently introduced simulation-based framework [6] to support the learning of non-parametric joint priors. The overall workflow remains unchanged; the modifications primarily concern how the prior distributions are specified and learned. We will first provide a brief overview of the workflow to establish the background, followed by a more detailed discussion of the modification introduced to the original framework, which is the focus of the current paper.
The general workflow of the framework closely resembles the approach of prior predictive checks [34]: We simulate from the joint model $p(\theta,y)$ and assess how well the resulting prior predictions align with the expert’s expectations. If there is a discrepancy between the expert’s expectations and the model simulations, the prior specification needs to be adjusted accordingly. Figure 1 provides a graphical representation of the framework and Table 1 offers a summary of the symbols and notation used in the following sections. The general workflow of our framework can be summarized as follows:
@@ 2.1 Prior Specification & Learning
In our previous work [6], we introduced the simulation-based framework along with a method for learning a set of independent parametric prior distributions, $p(\theta_{k}\mid\lambda_{k})$, for the model parameters $\theta_{k}$ with $k=1,\ldots,K$. In this paper, we extend our framework by introducing a method to learn a joint non-parametric prior distribution $p_{\lambda}(\theta)=p_{\lambda}(\theta_{1},\ldots,\theta_{K})$ over all model parameters $\theta$. In both the parametric and non-parametric approaches, the objective of the optimization process is to learn the hyperparameters $\lambda$ that minimizes the discrepancy between the model simulations and the expert expectations. However, the interpretation of these hyperparameters differs between approaches, which we emphasize by using a different notation for the prior distributions. In the non-parameteric approach, we employ normalizing flows (NFs) to induce a flexible family of prior distributions which entail specialized deep neural networks with trainable parameters $\lambda$ [46]. Thus, in the non-parametric approach, $\lambda$ represents the weights of the deep neural networks within the NFs, whereas in the parametric approach, $\lambda$ denotes the parameters of the associated prior distribution families (i.e., the prior hyperparameters). A more detailed introduction to NFs will be provided in the upcoming section.
In summary, the main difference between the parametric approach (proposed by [6]) and the non-parametric approach (focus of the current paper) is that, in the former, the user must specify a parametric prior distribution family for each model parameter, and the corresponding prior hyperparameters are learned via stochastic optimization. In the new extension, the user no longer needs to specify a parametric prior for the model parameters. Instead, a complex non-parametric prior distribution is assumed. This non-parametric joint prior is represented and learned using deep neural networks (specifically, NFs), with the weights learned via stochastic optimization. Table 2 summarizes the main steps in the elicitation process and highlights where the extension modifies the existing workflow of our simulation-based framework. Furthermore, we provide in this paper additional diagnostics to support users in interpreting the simulation results.
Elicitation process parametric-prior approach deep-prior approach
Setup
Model specify generative model %
Priors specify parametric prior distribution families assume non-parametric joint prior (specified via NFs)
Data specify target quantities and elicited statistics %
(less expert input needed due to stronger regularized priors) (more expert input required (e.g., dependency information) due to flexible prior)
Elicitation specify elicitation techniques used to query the domain expert %
Fitting apply mini-batch stochastic gradient descent %
Evaluation provide diagnostics, visualizations, etc. to evaluate simulation results %
Table 2: Comparison between the deep-prior and parametric-prior approach on a conceptual level. Note: The symbol ‘%’ refers to the same content as provided in parametric-prior approach. NFs is an abbreviation for Normalizing Flows (see text for details).
@@
NFs transform a simple probability distribution, known as the base distribution $p(u)$, into a more complex target distribution, which, in our case, is the joint prior $p_{\lambda}(\theta)$. This transformation is achieved through a series of invertible and differentiable mappings $g_{\lambda}$, parameterized by $\lambda$ [46]. The composition of invertible functions $g_{\lambda}$ yields an explicit form of the density $p_{\lambda}(\theta)$ via the change of variables formula
@@
Typically, $p(u)$ is chosen to be a simple density, such as a standard Gaussian or a uniform distribution [46].
@@ 2.2 Target Quantities & Elicitation Techniques
In our elicitation method, we represent an elicitation technique as a function $f_{j}$ for $j=1,\ldots,J$, of a target quantity $z_{i}$. The resulting set of elicited target quantities is referred to as elicited statistics,  $\{t_{m}(\lambda)\}$, where $t_{m}(\lambda)=f_{j}(z_{i})$ and $m$ refers to the corresponding $i\times j$ combination. In the following, we sometimes omit the explicit dependence on $\lambda$ and simply write $t_{m}$ instead of $t_{m}(\lambda)$. Note also that, depending on the elicitation technique employed, $t_{m}$ may accept either a single value or a set of values as input. To illustrate this point more clearly, consider the case in which we elicit expert knowledge about the predictive distribution conditional on a category of a categorical predictor, i.e., $z=y\mid\text{cat}$. To elicit this information, we query the expert regarding the $25\%,50\%$, and $75\%$ quantiles of this predictive distribution. Thus, we have $t=\{Q_{25\%}(y\mid\text{cat}),\,Q_{50\%}(y\mid\text{cat}),\,Q_{75\%}(y\mid\text{cat})\}$.
@@ 2.3 Model-implied & Expert-elicited Statistics
After determining the set of target quantities and corresponding elicitation techniques, we can proceed to gather information from the expert through specific prior elicitation protocols [37, 12, 23]. This process may be preceded by a training phase to familiarize the expert with the elicitation tasks [71]. In prior elicitation research, an entire subfield has focused on determining how to conduct optimal expert interviews, taking into account factors such as cognitive biases and heuristics. While it is beyond the scope of this paper to review this literature, several comprehensive reviews are available [e.g., 52, 71, 24]. At this point, we will skip the actual interrogation process and assume that the expert information has been successfully gathered, providing us with a set of expert-elicited statistics, individually denoted by $\hat{t}_{m}$ and together as $\{\hat{t}_{m}\}$.
@@ 2.4 Discrepancy Evaluation & Training
where $\alpha_{m}$ denote the weights. The choice of weights $\alpha_{m}$ can be guided by different considerations: One approach involves manually adjusting the weights, guided either by the expert’s assessment of the importance of individual loss components or by ensuring that the effects of different magnitudes, arising from varying scales, are balanced, such that no single loss component dominates the others [74]. Alternatively, automated weighting methods, known as loss-balancing methods, can be used to optimally balance the contribution of each individual term to the total gradient [48, 13, 3].
@@ 2.5 Evaluation of Training Results
Once the elicited statistics have been learned accurately, we can examine the corresponding learned prior distribution. As there are often multiple optimal values $\lambda^{*}$ that align closely with the same set of elicited statistics, multiple replications will yield different prior distributions [69, 49, 71]. This lack of uniqueness is anticipated, given the limited information provided by the set of elicited statistics compared to the complexity of the generative model [49]. Furthermore, when target quantities refer solely to the observable space, we only gain access to the model’s overall uncertainty with limited ability to differentiate between aleatoric and epistemic uncertainty [55, 62]. Aleatoric uncertainty, by definition, is irreducible; however, several methods exist to reduce epistemic uncertainty. These include (i) selecting more informative target quantities, (ii) eliciting more detailed expert information (e.g., increasing the number of quantiles in quantile-based elicitation), (iii) adding theory-informed regularization terms to constrain the search space [as illustrated in 49], (iv) improving the initialization method to achieve better starting points, and (v) employing more advanced optimization algorithms when dealing with a large number of model parameters [55].
@@
This oracle approach allows us to assess the self-consistency of the algorithm, the degree of non-identifiability (i.e., how informative are the elicited statistics), and the suitability of the chosen algorithm parameters used for training NFs with mini-batch SGD (such as the learning rate or the number of epochs). However, there is one important caveat when using an oracle: the user might be tempted to compare the learned prior distributions against the true prior distributions (which are now available). Specifically, if there is an unsatisfactory match between the true and learned prior distributions, the user might interpret this mismatch as a failure in learning. It is important to remember, however, that the method only knows about the elicited statistics, which are incorporated into the loss function. Thus, when the elicited statistics do not incorporate information about the model parameters but instead pertain to observable quantities, there is, from the method’s perspective, no concept of a true prior. In particular, if the true and learned elicited statistics match but there is a mismatch in the learned prior, this is likely a result of an underidentified problem, and the objective function requires additional information to properly constrain the problem.
@@ 3 Implementation
In all simulation studies, we use a batch size of $128$. To simulate from an oracle (discussed in Section 4.1), we draw once $S=10,000$ samples from the true joint prior and then compute the corresponding set of “expert”-elicited statistics. During training with mini-batch SGD we draw $S=200$ samples from the joint prior. To evaluate the discrepancy between the model-implied and expert-elicited statistics, we use a squared, biased Maximum Mean Discrepancy (MMD) loss with an energy kernel [38, 27], following our previous success with this metric [6]. For the correlation between parameters, however, we apply a squared-error loss (L2 loss; for details see Section 4.1). The total loss is computed as a weighted sum of the individual loss components. The values of the respective weights are discussed together with the selection of the elicited statistics whereby each elicited statistic represents one loss component (see Section 4.1).
The architecture of the NFs is rather simple compared to typical applications in deep learning [e.g., 44] and consists of a standard multivariate Gaussian as a base distribution and three affine coupling blocks. Each coupling block consists of two dense layers with $128$ units and ReLU activation functions. In Simulation Study 1, the number of learnable hyperparameters is $|\lambda|=202,776$, while in the case studies corresponding to Simulation Study 2, it is $|\lambda|=205,872$.
In Simulation Study 1, we assume a discrete likelihood and use the softmax-gumble trick to compute the gradients of a discrete random variable [for more information, see 6]. The gradients of the total loss with respect to the hyperparameters are used to update the hyperparameters $\lambda$. Optimization was performed with the Adam optimizer for $500$ epochs in Simulation Study 1 and $800$ epochs in Simulation Study 2. We used a fixed learning rate of $0.00025$ for Simulation Study 2, Scenario 1 & 2, and $0.0001$ for Simulation Study 1 and Stimulation Study 2, Scenario 3. These algorithm (hyper-)parameter values were obtained by manual tuning until good performance across all of our case studies was achieved.
All simulations and plots were performed in Python 3.11 using our Python package elicito 0.3.1 [5], which relies, among others, on the following libraries: TensorFlow 2.14 [1], TensorFlow Probability 0.23 [18], NumPy 1.26 [39], pandas 2.2.3 [51], and BayesFlow 1.1.6 [65]. To ensure the reproducibility of our results, the specific version of elicito used in this paper has been archived on Zenodo and is accessible via the following DOI https://doi.org/10.5281/zenodo.15241853. Simulations were executed on a YOGA Pro 9i laptop equipped with GPU acceleration.
@@ 4.1 General Setup
In the following, we present four case studies organized across two simulation studies to evaluate the performance of our elicitation method. The selection of case studies is motivated by the desire to choose scenarios that are simple and allow us to study the behavior of the method in a well-controlled setting. Since this paper presents the extension with NFs for the first time, it is important for us to first develop an understanding of the method’s behavior. In future work, we will examine the behavior of our method in more complex models that are closer to real-world applications.
In Simulation Study 1, we introduce a simple Binomial regression model with one continuous predictor. In Simulation Study 2, we present three scenarios based on a normal regression model with a three-level categorical predictor. The three normal scenarios differ only in the pre-defined joint prior representing the ground truth. Table 3 provides an overview of the generative models and the joint priors representing the ground truth used in the simulation studies. Before introducing the simulation studies, we will discuss the decisions made in the simulation setup in greater detail. All code and results can be found on GitHub (https://github.com/florence-bockting/non-parametric-prior) and OSF (https://osf.io/xrzh6/), respectively. Information about the training time for each simulation study is provided in Appendix A.1.
@@
Table 3: Overview of generative models and ground truth used in each simulation study. Symbols: $\mathbf{R}$ denotes the correlation matrix and $s$ the vector of standard deviations.
@@
For the simulation studies in this paper, we simulate expert input using synthetic data generated from a predefined joint prior representing the ground truth (i.e., oracle, see Section 2.5 for details). For simplicity, we use a parametric prior with fixed hyperparameter values as ground truth for each case study. Further details on the exact specification of the “true” joint prior is provided in the respective section and in Table 3.
@@
Following recommendations in the field, we tried to select mainly observable quantities as target quantities for our simulation studies (see Section 2.2 for details). For the Binomial regression model, we used the expected observations of the outcome variable conditional on two observations of the continuous predictor, $y\mid x_{0}$ and $y\mid x_{1}$. For the normal regression model, we considered the prior predictive observations conditional on three groups of the categorical predictor, $y\mid\text{gr}_{i}$ for $i=1,2,3$.
@@
Regarding the selection of elicitation techniques, we follow the recommendation to query quantiles from a domain expert [42]. For each target quantity, except for the correlation information, we elicit five quantiles: $5\%,25\%,50\%,75\%,95\%$. For the correlation values, we assume that the expert provides a single point estimate corresponding to a moment-based elicitation approach.
In summary, the elicited statistics for Simulation Study 1 consist of five quantiles each for $y\mid x_{0}$ ($t_{1}$) and $y\mid x_{1}$ ($t_{2}$), as well as the correlation information $\rho(\beta_{0},\beta_{1})$ ($t_{3}$). Each of these three elicited statistics corresponds to a distinct loss component in the objective function. The elicited statistics for Simulation Study 2 consist of five quantiles each for $R^{2}$ ($t_{1}$) and $y\mid\text{gr}_{i}$ with $i=1,2,3$ ($t_{2},t_{3},t_{4}$), as well as the correlation information between the model parameters ($t_{5}$). The discrepancy in the correlation values is measured using an L2 loss, whereas the discrepancies for the remaining components are computed using the $\text{MMD}^{2}$ loss. To account for differences in scale—arising both from the use of different discrepancy measures and from the varying domains of the elicited statistics—we weight the individual loss components as follows: the correlation loss component is scaled by a factor of $0.1$, the $R^{2}$ loss component by a factor of $10.0$, and all other loss components are uniformly weighted with a factor of $1.0$.
To assess whether the selected set of elicited statistics is sufficiently informative to determine a prior distribution for the model parameters, we conduct a sensitivity analysis for each case study.
@@
We run each simulation study 30 times with different random seeds. As preliminary convergence checks, we calculate the slope of the loss trajectory over the last 100 epochs and perform a visual convergence check for the five replications with the highest absolute slopes (i.e., worst cases). Additionally, we examine the trajectories of the total loss, the individual loss components related to each elicited statistic, as well as the mean and standard deviation for each marginal prior across epochs. Furthermore, to visually assess whether the learning criterion has been accurately captured, we compare the final model-implied statistics with the expert-elicited statistics. Finally, we examine the learned prior distributions for each replication and compute the model average across the 30 learned prior distributions. Table 4 provides a summary of the analysis workflow followed in each of the subsequent simulation studies.
@@
Additional simulation results
For each case study, we conduct two additional sets of simulations, which are not part of the main analysis but are included in the supplementary material for completeness and further comparison. The supplementary material can be found on GitHub at https://github.com/florence-bockting/non-parametric-prior. As these simulations are complementary to the main focus of the paper, we do not report the results in the main text.
In the first additional set of simulations, we re-ran all case studies using the model parameters as the “elicited statistics” (instead of an indirect quantity derived from the model parameters). In this case, the method trains directly on the model parameters, enabling us to assess whether it is, in principle, capable of recovering the true joint prior distribution when complete information is available. Thus, this set of simulations serves as a validation set for our method.
In the second additional set of simulations, we applied the parametric-prior method proposed in Bockting et al [6] to all case studies. This allows for a direct comparison between the performance of the proposed non-parametric approach and the previously introduced parametric variant within the same simulation-based framework. The only exception is Scenario 3 of Simulation Study 2, where we introduce a multivariate normal prior. In this case, a comparison is currently not possible, as the optimization of a full covariance matrix is not yet implemented in elicito [5], but it is planned for a (near) future release.